OpenStudy (anonymous):

simplify 2 root 50- 3 root 8

OpenStudy (saifoo.khan):

\[2\sqrt{50}-3\sqrt8\]?

OpenStudy (anonymous):

yea simply

OpenStudy (saifoo.khan):

Can you simplify sqrt(50) ?

OpenStudy (anonymous):

yea 2*25 and 5*10

OpenStudy (saifoo.khan):

Cool. Now: sqrt(50) = sqrt(2*25) = 5 sqrt2 What about 3sqrt8?

OpenStudy (anonymous):

3 cant but 2*4 is 8 @saifoo.khan

OpenStudy (saifoo.khan):

Right. So: sqrt(8) simplifies to: sqrt(2*4) = 2sqrt2 Can you solve the sum now?

OpenStudy (anonymous):

5 root 2 - 2 root 2 @saifoo.khan

OpenStudy (saifoo.khan):

You are forgetting the 2 and the 3 which were there in the start of the problem.

OpenStudy (anonymous):

do i add them to the problem i wrote?

OpenStudy (saifoo.khan):

2 root 50- 3 root 8 2*5 root 2 - 3*2 root 2